OpenStudy (anonymous):

What are the solutions of the system? y=x^2+4x-5 y=2x-2

OpenStudy (jit4won):

y=x^2+4x-5 y=2x-2 so 2x-2 =x^2+4x-5 you will get X^2 + 2x - 3 = 0

OpenStudy (jit4won):

(X + 3) ( x - 1)

OpenStudy (anonymous):

y=x^2+4x-5 y= 2x-2 SUBTRATCTING, 0=x^2+2x-5

OpenStudy (jit4won):

so x = - 3 or 1

OpenStudy (jit4won):

@ashwinjohn3 You made a mistake

OpenStudy (anonymous):

where???

OpenStudy (jit4won):

y=x^2+4x-5 y= 2x-2 SUBTRATCTING, 0=x^2+2x-3

OpenStudy (anonymous):

@jit4won thanks...alot

OpenStudy (anonymous):

x=-3,1

OpenStudy (anonymous):

oops. @jit4won just said that :]

OpenStudy (jit4won):

Yeah ...Unfortunately I am too quick

OpenStudy (anonymous):

canu guys do this question http://openstudy.com/study#/updates/4fa8c344e4b059b524f4b6c0?